OpenStudy (anonymous):

someone please help me. Verify the identity. Justify each step. tan θ + cot θ = 1/ sin θcos θ

OpenStudy (anonymous):

\[\tan x+\cot x=\frac{ \sin x }{ \cos x }+\frac{ \cos x }{ \sin x }\] \[=(\frac{ \sin x }{ \cos x })(\frac{ \sin x }{ \sin x }+(\frac{ sinx }{ \cos x })(\frac{ \cos x }{ \cos x })\] \[=(\frac{ (\sin x)^{2} }{ sinx \cos x })+(\frac{ (\cos x)^{2} }{ sinx \cos x })\] \[=\frac{ ((sinx)^{2}+(\cos x)^{2} )}{ sinx \cos x}\] \[=\frac{ 1 }{ sinx cosx }\]

OpenStudy (anonymous):

thanks :)